### v1

# Long-term Causal Inference via Modeling Sequential Latent Confounding

Abstract
Long-term causal inference is an important but challenging problem across various scientific domains. To solve the latent confounding problem in long-term observational studies, existing methods leverage short-term experimental data. Ghassami et al. (2022) propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified. While effective in certain cases, this assumption is limited to scenarios with a one-dimensional short-term outcome. In this paper, we introduce a novel assumption that extends the CAECB assumption to accommodate temporal short-term outcomes. Our proposed assumption states a functional relationship between sequential confounding biases across temporal short-term outcomes, under which we theoretically establish the identification of long-term causal effects. Based on the identification result, we develop an estimator and conduct a theoretical analysis of its asymptotic properties. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
1 Introduction
Long-term causal inference is an important but challenging problem across various scientific fields, such as education (Athey et al., 2019), medicine (Fleming et al., 1994), and marketing (Hohnhold et al., 2015). While in many real-world scenarios, long-term observational data are readily available, a major challenge for long-term causal inference is the presence of latent confounding in observational studies. A common way to mitigate this issue is incorporating short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Figure 1?
Figure 1: Causal graphs of experimental data and observational data with $\mathbf{X}$ being covariates, $\mathbf{U}$ being latent confounders, $A$ being treatment, $\mathbf{S}$ being short-term outcome, and $Y$ being long-term outcome. Fig. 1(a) represents the causal graph of the short-term experimental data, where treatment $A$ is not affected by latent confounders $\mathbf{U}$ and the long-term outcome $Y$ is unobserved. Fig. 1(b) represents the causal graph of the long-term observational data, where the latent confounders $\mathbf{U}$ affect treatments $A$ and outcomes $\mathbf{S},Y$ and the long-term outcome $Y$ can be observed. Fig. 1(c) represents the overall graph with expanding temporal short-term outcomes.
Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions. One widely used assumption is the Latent Unconfoundedness assumption (Athey et al., 2020; Chen & Ritzwoller, 2023; Yang et al., 2024), which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A|\mathbf{S}(a),\mathbf{X},G=O$. These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States (Aizer et al., 2024). However, this assumption essentially restricts that the latent confounders can not affect the long-term outcome, i.e., ruling out the causal edge $\mathbf{U}\rightarrow Y$. To address this limitation, a follow-up work (Ghassami et al., 2022) proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$. The CAECB assumption states that the confounding bias in the short-term outcome is equal to that in the long-term outcome, enabling the identification of long-term causal effects. This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes (Jackson, 2016).
Following the work of Ghassami et al. (2022), we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists. Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption (Ghassami et al., 2022) is not applicable. In many real-world applications, short-term outcomes exhibit temporal dependencies, and capturing these relationships is essential for inferring long-term causal effects. For example, in evaluating the long-term (e.g. year-long) effectiveness of the medication, patients undergo regular follow-up visits (e.g., weekly or monthly). During these visits, temporal short-term health indicators are recorded as short-term outcomes. These sequential measurements capture the progression of patients’ conditions over time, providing valuable information on the long-term outcome of interest. Consequently, the absence of temporal considerations in existing methods limits their effectiveness and constrains the potential of long-term causal inference in practice.
In this paper, we introduce a novel assumption called Functional CAECB (FCAECB), which extends the existing CAECB assumption to capture the temporal dependencies among temporal short-term outcomes. Roughly speaking, the proposed FCAECB assumption posits a functional relationship between sequential latent confounding biases. Under this assumption, we establish the theoretical identification of long-term causal effects. Correspondingly, we devise an algorithm for the long-term effect estimation. Additionally, we analyze the convergence rates of our proposed estimator in the asymptotic and finite sample setting within a generic nonparametric regression framework, with the ultimate goal of deepening the understanding of how sequential short-term confounding biases contribute to inferring long-term effects. Overall, our contribution can be summarized as follows:
  • Novel Assumption for Long-Term Causal Inference: We study the problem of long-term causal inference in the presence of temporal short-term outcomes. We propose a novel assumption named FCAECB for identifying long-term causal effects, which enables capturing the time-dependent relationships between sequential latent confounding biases. Note that the existing CAECB assumption can be seen as our special case.
  • Estimator for Heterogeneous Long-Term Effects: We devise an estimator for estimating heterogeneous long-term effects under the proposed FCAECB assumption, which can be implemented using any machine learning regression method. Theoretically, we analyze the asymptotic properties of our estimator.
  • Empirical Validation: We conduct extensive experiments to validate our theoretical findings and demonstrate the effectiveness of the proposed estimator.
2 Related Work
Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects. Various criteria have been proposed for identifying valid surrogates, including the Prentice criteria (Prentice, 1989), principal surrogacy (Frangakis & Rubin, 2002), strong surrogate criteria (Lauritzen et al., 2004), causal effect predictiveness (Gilbert & Hudgens, 2008), and consistent surrogate and its variations (Chen et al., 2007; Ju & Geng, 2010; Yin et al., 2020). Recently, there has been growing interest in estimating long-term causal effects using surrogates, which is also the focus of this paper. One line of work assumes the unconfoundedness assumption. Under the unconfoundedness assumption, LTEE (Cheng et al., 2021) and Laser (Cai et al., 2024) are based on specifically designed neural networks for long-term causal inference. EETE (Kallus & Mao, 2020) explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effect. ORL (Tran et al., 2023) introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes. Singh & Sun (2024) proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment. Additionally, Wu et al. (2024) develop a policy learning method for balancing short-term and long-term rewards. Our work is different from them. We do not assume the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders. Another line of research, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data. This setting is first introduced by the method proposed by Athey et al. (2019), which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification. As follow-up work, Athey et al. (2020) introduces the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification. Under this assumption, Yang et al. (2024); Chen & Ritzwoller (2023) propose several estimators for effect estimation. The alternative feasible assumptions (Ghassami et al., 2022) are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption. Based on proximal methods (Tchetgen Tchetgen et al., 2024), Imbens et al. (2024) proposes considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification. However, these studies primarily concentrate on average treatment effects, whereas our focus in this paper is on heterogeneous effects. Among the existing literature, the most closely related work is Ghassami et al. (2022)’s work. Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of the assumption we propose.
Modeling Latent Confounding Bias An effective way to solve the latent confounding problem in the data combination setting is to model latent confounding bias. Kallus et al. (2018) proposes modeling confounding bias under a linearity assumption. Hatt et al. (2022) introduce to model the nonlinear confounding bias using the representation learning technique. Wu & Yang (2022) propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality. Zhou et al. (2025) propose a two-stage representation learning strategy to model such a confounding bias. Different from these works, we focus on the long-term causal inference setting, and rather than focus on how to model the confounding bias, we concentrate more on the relationship between sequential confounding biases.
3 Preliminaries
3.1 Notations, Problem Definition, Assumptions
Let $A\in\{0,1\}$ be the treatment variable, $\mathbf{X}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be the observed covariates where $d$ is the dimension of $\mathbf{X}$, $\mathbf{U}\in{\mathcal{U}}\subseteq{\mathbb{R}}^{d_{U}}$ be the latent confounders where $d_{U}$ is the dimension of $\mathbf{U}$. Let $\mathbf{S}=[S_{1},S_{2},\dots,S_{T}]$ be the short-term outcome variable where $S_{t}\in{\mathcal{S}}\subseteq{\mathbb{R}}$ is the short-term outcome measured at timestep $t$, and $Y=S_{T+\mu}\in{\mathcal{R}}\subseteq{\mathbb{R}}$ be the long-term outcome. Further, we leverage the potential outcome framework proposed by Rubin (1978). We denote $\mathbf{S}(a)$ as the potential short-term outcome, $S_{t}(a)$ as the potential short-term outcome at timestep $t$, and $Y(a)$ as the potential long-term outcome. Following existing work on long-term inference (Athey et al., 2020; Ghassami et al., 2022; Imbens et al., 2024; Chen & Ritzwoller, 2023), we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data. Let lowercase letters (e.g., $a,\mathbf{x},\mathbf{u},\mathbf{s},y,s(a),y(a)$) denote the value of the previously described random variables. Let the superscript $(i)$ denote a specific unit, e.g., $\mathbf{x}^{(i)}$ is the covariate value of unit $i$. Then, the experimental data and the observational data are denoted as $\mathbb{D}_{e}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and $\mathbb{D}_{o}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},G^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, where $n_{e}$ and $n_{o}$ are the size of experimental data and the observational data respectively.
For ease of convenience, we denote the following nuisance functions and confounding bias:
$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=E],$ (1)
$\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=O],$
$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=E],$
$\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=O],$
$\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X}),$
where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
Moreover, we denote stochastic boundedness with $O_{p}$ and convergence in probability with $o_{p}$. We denote $X_{1}{\perp\!\!\!\perp}X_{2}$ as the independence between $X_{1}$ and $X_{2}$. We use $a_{n}\asymp b_{n}$ to denote both $a_{n}/b_{n}$ and $b_{n}/a_{n}$ are bounded. We use $a_{n}\lesssim b_{n}$ to denote both $a_{n}\leq Cb_{n}$ for some constant $C>0$.
Task: Given a short-term experimental dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and a long-term observational dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},g^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, the goal is to identify the following causal estimand of interest: the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}].$ (2)
However, HLCE $\tau(\mathbf{x})$ is not identifiable from the experimental data alone, since $Y$ is missing in that dataset. Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem. Furthermore, the information regarding the causal effects in the experimental data is not necessarily relevant to that in the observational data without further assumptions. To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference:
Assumption 3.1 (Consistency).
If a unit is assigned treatment, we observe its associated potential outcome. Formally, if $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
Assumption 3.2 (Positivity).
The treatment assignment is non-deterministic. Formally, $\forall a,\mathbf{x}$, we have $10<P(A=a|\mathbf{X}=\mathbf{x})<$ 1, and $0<P(G=O|A=a,\mathbf{X}=\mathbf{x})<1$.
Assumption 3.3 (Weak internal validity of observational data).
Latent confounders exist in observational data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},\mathbf{U},G=O$ and $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=O$.
Assumption 3.4 (Internal validity of experimental data).
There are no latent confounders in experimental data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=E$.
Assumption 3.5 (External validity of experimental data).
The distribution of the potential outcomes is invariant to whether the data belongs to the experimental or observational data. Formally, afor-$\forall a$, we have $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X}$.
Assumptions 3.1 and 3.2 are standard assumptions in causal inference (Rubin, 1974; Imbens, 2000). Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings (Shi et al., 2023; Imbens et al., 2024; Athey et al., 2019, 2020; Hu et al., ). Specifically, Assumption 3.3 allows the existence of latent confounders in observational data, thus it is much weaker than the traditional unconfoundedness assumption. Assumption 3.4 is reasonable and can hold since the treatment assignment mechanism is under control in the experiments. Assumption 3.5 connects the potential outcome distributions between observational and experimental data.
Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest. The root cause is that, even though the assumptions above link the experimental and observational data, the (long-term) latent confounding problem remains unsolved. In the following section, we first review a method proposed by Ghassami et al. (2022), which poses an extra assumption, called Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE. Then, we propose our approach that generalizes the method of Ghassami et al. (2022) to allow temporal short-term outcomes.
(a) Schematic representation of CAECB assumption (Ghassami et al., 2022).
(b) Schematic representation of our proposed FCAECB assumption.
Figure 2: Schematic representations of CAECB assumption proposed by Ghassami et al. (2022) and our FCAECB assumption. As shown in Figure 2(a), the CAECB assumption requires that the confounding bias in the short-term outcome is equal to that in the long-term outcome. As shown in Figure 2(b), the FCAECB assumption relaxes this constraint by allowing for temporal short-term outcomes and only requiring that confounding biases across different time steps follow a specific pattern rather than remaining equal.
3.2 CAECB Assumption (Ghassami et al., 2022)
Ghassami et al. (2022) introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption. The method under CAECB assumption is only applicable to scenarios with a one-dimensional short-term outcome, which we denote as $S$ (and its corresponding potential outcome $S(a)$) with slightly abusing notation.
To begin with, we first restate the CAECB assumption as follows:
Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB (Ghassami et al., 2022)).
The difference in conditional expected values of short-term potential outcomes across treated and control groups is the same as that of the long-term potential outcome variable. Formally, afor-$\forall a$, we have
[…]
Theorem 3.7.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (4)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$
where $\omega(\mathbf{x})=\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x})$ is the short-term confounding bias.
Proof can be found in Appendix A. A similar identification result in terms of long-term average causal effects has been shown by Ghassami et al. (2022). Here we provide the identification result of HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
Remark 3.8.
We illustrate Assumption 3.6 in Figure 2(a). Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework (Ashenfelter & Card, 1984; Angrist & Pischke, 2009), Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one. In this way, $\omega(\mathbf{x})$ can be interpreted as the long-term confounding bias, resulting in the identification result in Eq. (4).
Considering Eq. (3) in Assumption 3.6, $S$ and $Y$ should be on the same scale, which restricts the practical application of the method under this assumption. In the next section, we extend this assumption to allow for temporal short-term outcomes, which significantly improve the practical utility.
4 Long-term Identification under Functional CAECB Assumption
In this section, we introduce a novel assumption considering the temporal information between temporal short-term outcomes, enabling the identification of long-term causal effects.
Specifically, we formalize our proposed assumption:
Assumption 4.1 (Functional Conditional Additive Equi-Confounding Bias, FCAECB).
The difference in conditional expected values of short-term potential outcomes across treated and control groups between two timesteps follows the learnable function forms $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$. Formally, afor-$\forall a$, we have
$b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X}),$ (5)
where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)|\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)|\mathbf{X},A=1,G=O]$.
Remark 4.2.
We illustrate our proposed FCAECB assumption in Figure 2(b). We relax the CAECB assumption (Ghassami et al., 2022) to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome. Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCEACB assumption satisfies $f(\mathbf{X})=1$.
To provide a better understanding of our Assumption 4.1, we provide the insight in term of the functional form $f(\mathbf{X})$ in the following proposition.
Proposition 4.3.
Under Assumption 4.1, $\forall t$ the confounding biases between times $t$ and $t+1$ follow
$\omega_{t+1}(\mathbf{X})=f(\mathbf{X})\omega_{t}(\mathbf{X}),$ (6)
where $\omega_{t}(\mathbf{X})$ is the confounding bias at time step $t$, defined as $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$.
The proof is given in Appendix B.
Remark 4.4.
As stated in Proposition 4.3, Assumption 4.1 essentially states the confounding biases $\omega_{t}(\mathbf{X})$ between adjacent time steps follow the functional form $f(\mathbf{X})$. Proposition 4.3 also illustrates the way of how to learn the function $f(\mathbf{X})$ using the observed variables, unlike the definition in Eq. (5) defining $f(\mathbf{X})$ using the potential outcomes (See Section 5 for the estimation based on Proposition 4.3).
The key to identifying long-term causal effect under Assumption 4.1 is using the temporal information in $f(\mathbf{X})$ to extrapolate the long-term confounding bias $\omega_{T+\mu}(\mathbf{X})$. Using Eq. (6), the long-term confounding bias $\omega_{T+\mu}(\mathbf{X})$ can be expressed as $\omega_{T+\mu}(\mathbf{X})=f^{\mu}(\mathbf{X})\omega_{T}(\mathbf{X})$, which results in the following theorem of long-term effect identification.
Theorem 4.5.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then the heterogeneous long-term effects $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (7)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$
where $\omega_{T}(\mathbf{x})=\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$ is the short-term confounding bias at time step $T$.
The proof is given in Appendix C. Theorem 4.5 provides the identification result of the heterogeneous long-term effects $\tau(\mathbf{X})$. The identification result consists of two parts: 1. long-term outcome differences in observational data $\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and 2. long-term confounding bias $f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x})$. The long-term confounding bias is identified by the extrapolated result using short-term confounding bias $\omega_{T}(\mathbf{x})$ under Assumption 4.1. The identification result also illustrates a way to estimate the long-term effects $\tau(\mathbf{x})$ via modeling sequential latent confounding, which inspires our estimator as shown in Section 5.
5 Long-term Causal Effect Estimation
In this section, we first introduce our estimator $\hat{\tau}(\mathbf{X})$ for heterogeneous long-term effects under our proposed Assumption 4.1, and provide the corresponding theoretical analysis of the proposed estimator.
5.1 Estimator
Our estimator $\hat{\tau}(\mathbf{X})$ directly follows the identification result in Theorem 4.5. As shown in Eq. (7), the estimators consist of three nuisance components, outcome mean difference between treated and control group in observational data $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})$, confounding bias $\omega_{T}(\mathbf{x})$, and the function between confounding bias $f(\mathbf{x})$. The first and the second terms can be directly estimated by fitting nuisance functions, and the third term $f(\mathbf{x})$ should be estimated based on the fitted confounding biases. Accordingly, we design our heterogeneous long-term effect estimator within a two-stage regression framework, which are model-agnostic algorithms that decompose the estimation task into multiple sub-problems, each solvable using any supervised learning/regression methods.
Specifically, our proposed estimator follows:
  • S0.
    (Optional) Selecting subsets of short-term outcomes $\mathbf{S}$, yielding appropriate $T$ and $\mu$;
  • S1.
    Fitting the following nuisance functions: $\hat{\mu}_{Y}^{O}(a,\mathbf{x})$, $\hat{\mu}_{S_{t}}^{E}(a,\mathbf{x})$, and $\hat{\mu}_{S_{t}}^{O}(a,\mathbf{x})$ for all $a$ and $t$;
  • S2.
    Constructing the confounding bias $\hat{\omega}_{t}(\mathbf{x})=\hat{\mu}_{S_{t}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{t}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{t}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{t}}^{O}(1,\mathbf{x})$ for all $t$, and fitting the function $\hat{f}(\mathbf{x})$ by minimizing
    $arg\!\min_{f\in{\mathcal{F}}}\Sigma_{t=1}^{T-1}\Sigma_{i=1}^{n}\left(\hat{\omega}_{t+1}(\mathbf{x}_{i})-f(\mathbf{x}_{i})\hat{\omega}_{t}(\mathbf{x}_{i})\right)^{2};$ (8)
  • S3.
    Constructing final HLCE estimator as
    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})$ (9)
    ×$\times\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$
Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional S0 to select subsets of short-term outcomes $\mathbf{S}$. This step is motivated by the identification result in Eq. (7) that allows for different choices of $\mu$ and $T$. For example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. Then we have multiple choices of $T$ and $\mu$, e.g., using all short-term outcomes $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ with $T=6$ and $\mu=3$, or using $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3$ and $\mu=2$, or using $\mathbf{S}=[S_{1},S_{5}]$ with $T=2$ and $\mu=1$. We will discuss how to choose $T$ and $\mu$ in the next section.
5.2 Convergence Rate Analyses
In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate (Stone, 1980) of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$. Formally, we provide the following definition.
Definition 5.1 (Hölder ball).
The Hölder ball $\mathcal{H}_{d}(s)$ is the set of $s$-smooth functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ supported on $\mathcal{X}\subseteq\mathbb{R}^{d}$ that are $\lfloor s\rfloor$-times continuously differentiable with their multivariate partial derivatives up to order $\lfloor s\rfloor$ bounded, and for which
$|\frac{\partial^{m}f}{\partial^{m_{1}}\cdots\partial^{m_{d}}}(x)-\frac{\partial^{m}f}{\partial^{m_{1}}\cdots\partial^{m_{d}}}(x^{\prime})|\lesssim\|x-x^{\prime}\|^{s-\lfloor s\rfloor}_{2},$
$\forall x,x^{\prime}$ and $m=(m1,\cdots,m_{d})$ such that $\Sigma_{j=1}^{d}m_{j}=\lfloor s\rfloor$.
Assumption 5.2 (Smoothness Assumption).
We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone (1980)’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
We also assume the boundedness assumption of the nuisance functions.
Assumption 5.3 (Boundness Assumption).
We assume that the nuisance functions $\mu_{S_{t}}^{O}(a,x),\mu_{S_{t}}^{E}(a,x)$ and their estimates are bounded, i.e., $\forall t$, $|\mu_{S_{t}}^{O}(a,x)|<C_{1}$, $|\hat{\mu}_{S_{t}}^{O}(a,x)|<C_{2}$, $|\mu_{S_{t}}^{E}(a,x)|<C_{3}$ and $|\hat{\mu}_{S_{t}}^{E}(a,x)|<C_{4}$ hold for some $C_{1},C_{2},C_{3},C_{4}>0$.
We now state our main theoretical results. To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from (Kennedy, 2023), which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively. Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting (Kennedy, 2023) and later is adapted to several other methods (Curth & Van der Schaar, 2021; Frauen & Feuerriegel, ). Different from them, we use such a technique for the sequential latent confounding modeling, which is then adapted for the long-term effect estimation.
To begin with, we first provide the rate of $\hat{f}(\mathbf{x})$ as follows:
Lemma 5.4.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, and 5.2 hold, then we have
$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\right.$ (10)
$\left.+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$
The proof is given in Appendix D. Since $\hat{f}(\mathbf{x})$ is a time-series model, its effective sample is $(T-1)n$, thus it might achieve a faster rate if we observed a longer duration of short-term outcomes. Moreover, the rate of $\hat{f}(\mathbf{x})$ consist two part: the oracle rate $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}$, and the rate of fitting nuisance functions $(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}$. If the nuisance functions $\mu_{S_{t}}^{E}$ and $\mu_{S_{t}}^{O}$ is smooth enough such that $\min\{\alpha,\beta\}\geq\eta$, then $\hat{f}(\mathbf{x})$ will attain the oracle rate.
Based on Lemma 5.4, we provide the rate of $\hat{\tau}(\mathbf{x})$ in the following theorem.
Theorem 5.5.
Suppose Lemma 5.4 hold, then we have
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})$ (11)
$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}\right.$
$\left.+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
Proof can be found in Appendix E. Theorem 5.5 follows directly from Lemma 5.4. The rate of $\hat{\tau}(\mathbf{x})$ consist of several terms: 1. $n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$, 2. $n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$, 3. and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$. Theorem 5.5 suggests that the convergence rate is primarily influenced by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal parameters $T,\mu$, where $T$ represents the longest observed duration of the short-term outcomes $\mathbf{S}$ and $\mu$ represents the time horizon of the long-term outcome to be estimated. This implies that achieving a more accurate effect estimation requires observing short-term outcomes for as long as possible.
Remark 5.6 (Choosing $\mu$ and $T$).
Note that, the bound in Eq. (11) contains the term $\frac{\mu}{(T-1)^{\frac{\alpha}{2\alpha+d}}}$ (similar for $\beta$ and $\eta$). If we have prior knowledge of the smoothness of nuisance functions, i.e., knowing $\alpha$ ($\beta$ and $\eta$), then we can optimally select $\mu$ and $T$ to obtain a faster rate. For instance, if all estimated functions are sufficiently smooth, i.e., $\alpha,\beta,\eta\rightarrow\infty$, then $\mu$ and $T$ should be chosen to minimize $\frac{\mu}{\sqrt{T-1}}$, since $\frac{\alpha}{2\alpha+d},\frac{\beta}{2\beta+d},\frac{\eta}{2\eta+d}\rightarrow 1/2$. As a concrete example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. The optimal choice is to use $\mathbf{S}=[S_{1},S_{5}]$ only to estimate $\tau(\mathbf{x})$, since it result in a minimum $\frac{\mu}{\sqrt{T-1}}=1$ where $\mu=1$ and $T=2$. In practice, when no prior knowledge is available, it may be advisable to choose $\mu$ as small as possible, since a larger $\mu$-power in $\hat{f}^{\mu}(\mathbf{x})$ could lead to significant errors.
6 Experiments
In this section, we perform a series of experiments to evaluate the effectiveness and validity of our proposed estimator. We first provide the experimental setup in Section 6.1, and provide the experimental results and corresponding analyses in Section 6.2.
6.1 Experimental Setup
In this section, we describe the data generation process for the synthetic datasets, the implementation details, and the metrics used in the experiments.
Synthetic Datasets The data generation process is partly following (Kallus et al., 2018) such that we can obtain a specific form of confounding bias. Specifically, we first generate the treatments as follows: $A|G=O\sim\text{Bernoulli}(0.6)$ and $A|G=E\sim\text{Bernoulli}(0.4)$. Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
$(\mathbf{X},\mathbf{U})|A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
$(\mathbf{X},\mathbf{U})|A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{\mu}$ satisfying Assumption 4.1 as follows:
$S_{t}(1)=1+1.1\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(1)+\epsilon_{S_{t}},$ (13)
$S_{t}(0)=\mathbf{X}+\mathbf{U}+\sum_{k=0}k^{t}S_{k}(0)+\epsilon_{S_{t}},$
where $\epsilon_{S_{t}}$ are Gaussian noises. This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=t\times\mathbf{x}$. We perform several control experiments using this data generation process. The default values are: the observational data sample size $n_{o}=4000$, experimental data sample size $n_{e}=2000$, $T=6$, and $\mu=3$. In control experiments, we vary $\mu$ within $\{1,2,3,4,5\}$ and $T$ within $T\in\{4,5,6,7,8\}$. We also fix the ratio of $n_{O}:n_{E}=2:1$, and varying $n_{e}$ within $\{1000,2000,3000,4000,5000,10000\}$ (which corresponds to $n_{o}$ values of $\{2000,4000,6000,8000,10000,20000\}$).
Implementation We fit all nuisance functions using correctly specified regression. We denote our estimator as $\hat{\tau}$. For comparison, we evaluate our method against a correctly specified T-Learner (Künzel et al., 2019), denoted as $\hat{\tau}_{exp}$, which directly regresses unobserved $Y$ in the experimental data. However, this estimator $\hat{\tau}_{exp}$ is infeasible in practice since $Y$ is missing in the experimental data. This baseline serves as an idealized benchmark, and comparing our method against it highlights the effectiveness of our estimator across different control experiments.
Metrics As for heterogeneous effect estimation, we report Precision in the Estimation of Heterogeneous Effect (PEHE) $\varepsilon_{PEHE}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\tau(\mathbf{x}_{i})-\hat{\tau}(\mathbf{x}_{i}))^{2}}$. As for average long-term causal effect estimation, we report the absolute error $\varepsilon_{ATE}=|\Sigma_{i=1}^{n}\tau(\mathbf{x}_{i})-\Sigma_{i=1}^{n}\hat{\tau}(\mathbf{x}_{i}))|$. For all metrics, we report the mean values and standard errors by $100$ times running.
6.2 Experimental Results
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
Figure 3: Results of the experiments in terms of different choice of $\mu$ and $T$.
Optimal Choice of $T$ and $\mu$: We conduct experiments using different choice of $T$ and $\mu$. Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$: 1. $\mathbf{S}=[S_{1},S_{3}]$ with $T=2,\mu=3$; 2. $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3,\mu=2$; 3. $\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5}]$ with $T=5,\mu=3$; and 4. $\mathbf{S}=[S_{1},S_{5}]$ with $T=2,\mu=1$. The results, shown in Figure 3, are sorted by the values of $\frac{\mu}{\sqrt{T-1}}$, which $3,1.41,1.34$, and $1$, respectively. As shown in Figure 3, as $\frac{\mu}{\sqrt{T-1}}$ decrease, both the estimation errors $\varepsilon_{PEHE}$ and $\varepsilon_{ATE}$ decrease. This aligns with our theoretical findings in Theorem 5.5, as a smaller $\frac{\mu}{\sqrt{T-1}}$ leads to a faster rate. Interestingly, as shown in Figure 5(b), we found that when the $\frac{\mu}{\sqrt{T-1}}$ is equal to $1$, our estimator $\hat{\tau}_{(}\mathbf{x})$ achieves lower $\varepsilon_{ATE}$ than the idealized estimator $\hat{\tau}$. This may be attributed to the fact that our estimator $\hat{\tau}(\mathbf{x})$ leverages observational data, leading to higher data efficiency.
(a) Heterogeneous effect estimation error with fixed $T$ and varying $\mu$.
(b) Average effect estimation error with fixed $T$ and varying $\mu$.
(c) Heterogeneous effect estimation error with fixed $\mu$ and varying $T$.
(d) Average effect estimation error with fixed $\mu$ and varying $T$.
Figure 4: Results of control experiments with fixed $T$ and varying $\mu$ (Figures 4(a) and 4(b)), and with fixed $\mu$ and varying $T$ (Figures 4(c) and 4(d)).
Vary $T$ and $\mu$: We conduct control experiments with varying $T$ and varying $\mu$ separately. First, Figures 4(a) and 4(b)) show the results for the experiments with fixed $T$ and varying $\mu$. As expected, the estimation errors on both heterogeneous and average effect estimations increase as the long-term horizon $\mu$ grows. This is primarily due to the increasing estimation errors of $\hat{f}^{\mu}(\mathbf{x})$, as established in Lemma 5.4. Second, Figures 4(c) and 4(d) show the results for the experiments with fixed $\mu$ and varying $T$. As the observed duration $T$ increases, the estimation errors decrease rapidly. It is reasonable since the larger $T$ results in a faster rate of $\hat{\tau}(\mathbf{x})$. Both findings further support our Lemma 5.4 and Theorem 5.5 again. Moreover, we observe that, with smaller $\mu$, our estimator closely matches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, showing the effectiveness of our estimator.
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
Figure 5: Results of control experiments with varying sample sizes $n_{e}$ and $n_{o}$.
Sample sensitivity: We conduct control experiments by varying the sample sizes of both experimental and observational data. The results are presented in Figure 5. As expected, the estimator’s performance improves as the sample sizes $n_{o}$ and $n_{e}$ increase. Notably, when $n_{o}$ and $n_{e}$ become larger, our estimator $\hat{\tau}(\mathbf{x})$ closely approaches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, which again confirms the effectiveness of our estimator. Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{e}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
7 Conclusion
In this paper, we extend the existing CAECB assumption (Ghassami et al., 2022) for long-term causal inference. The original assumption is restricted to settings with a one-dimensional short-term outcome. To address this limitation, we introduce a more general assumption—the functional CAECB assumption—which accommodates temporal short-term outcomes. We theoretically establish the identification result of the heterogeneous long-term causal effect under our assumption and propose a corresponding estimator by modeling sequential latent confounding. Additionally, we provide a comprehensive theoretical analysis of the estimator. Experiments confirm both the validity of our theoretical results and the effectiveness of the proposed estimator.
Impact Statement
This paper presents work whose goal is to advance the field of Long-term causal inference in the presence of latent confounders. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 3.7
We first restate Theorem 3.7 as follows:
Theorem A.1.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})$ (14)
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x}).$
Proof.
$\tau(\mathbf{X})$ (15)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=1\right]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=1\right]\}\times p(A=1|\mathbf{X},G=O),$
where the first equality is based on Assumption 3.5 and the last equality is based on Assumption 3.6. Similarly, for short-term conditional causal effects, we have:
$\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]$ (16)
$=$ $\mathbb{E}[S|\mathbf{X},G=O,A=1]-\mathbb{E}[S|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[S(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
Then, combining Eq. (15) and (16), we have
$\tau(\mathbf{X})$ (17)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=E]-\mathbb{E}[S(0)|\mathbf{X},G=E]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S|\mathbf{X},G=E,A=1]-\mathbb{E}[S|\mathbf{X},G=E,A=0]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mu_{Y}^{O}(1,\mathbf{X})-\mu_{Y}^{O}(0,\mathbf{X})+\mu_{S}^{E}(1,\mathbf{X})-\mu_{S}^{E}(0,\mathbf{X})+\mu_{S}^{O}(0,\mathbf{X})-\mu_{S}^{O}(1,\mathbf{X}),$
where the second equality is based on Assumption 3.5 and the last equality is based on Assumption 3.4. This finishes our proofs. ∎
Appendix B Proof of Proposition 4.3
We first restate the proposition as follow:
Proposition B.1.
Under Assumption 4.1, $\forall t$ the confounding biases between times $t$ and $t+1$ follow
$\omega_{t+1}(\mathbf{X})=f(\mathbf{X})\omega_{t}(\mathbf{X}),$ (18)
where $\omega_{t}(\mathbf{X})$ is the confounding bias at time step $t$, defined as $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$.
Proof.
We start from the long-term causal effects at time step $t$ in experimental data $G=E$:
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (19)
=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
where the equality $(a)$ is based on Assumption 3.5. By rewriting the last equality above, we obtain:
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (20)
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\iff$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$=$ $b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\iff$ $\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$
$=$ $b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\overset{(a)}{\iff}$ $\omega_{t}(\mathbf{X})=b_{t}(1,\mathbf{X}))p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
where $(a)$ is based on the definition of $\omega_{t}(\mathbf{X})$. Then, similarly for time step $t+1$, we have
$\omega_{t+1}(\mathbf{X})$ $=b_{t+1}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t+1}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$ (21)
=$\overset{(a)}{=}f(\mathbf{X})\left(b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)\right)$
=$\overset{(b)}{=}f(\mathbf{X})\omega_{t}(\mathbf{X}),$
where $(a)$ is based on Assumption 4.1, i.e., $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X})$, and $(b)$ is based on Eq. (20) and finishes our proof. ∎
Appendix C Proof of Theorem 4.5
We first restate Theorem 4.5 as follows:
Theorem C.1.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})$ (22)
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right).$
Proof.
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (23)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+b_{T+\mu}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T+\mu}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)|\mathbf{X}$, and last equality is based on Assumption 4.1.
Further based on Assumption 4.1, we have $b_{T+\mu}(a,\mathbf{X})=f^{\mu}(\mathbf{X})b_{T}(a,\mathbf{X})$. Then, we rewrite the equality above as:
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (24)
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+f^{\mu}(\mathbf{X})b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-f^{\mu}(\mathbf{X})b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
Similarly, for short-term ITE at last time step $T$ we have:
$\mathbb{E}[S_{T}(1)|\mathbf{X}]-\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (25)
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$+b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
Then, combining Eq. 24 and 25, we have
$\tau(\mathbf{X})=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)|\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (26)
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=0]$
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right)$
where the second equality is based on Assumption 3.5. This finishes our proof. ∎
Appendix D Proof of Lemma 5.4
We first state the following lemma that is used in the proof of Lemma 5.4.
Lemma D.1.
suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1 hold, then $\forall t$,
$\hat{\omega}_{t}(\mathbf{x})-\omega_{t}(\mathbf{x})=O_{p}\left(r_{\mu_{S_{t}}^{E}}(n)+r_{\mu_{S_{t}}^{O}}(n)\right).$ (27)
where $r_{\circ}(n)$ denotes the risk of nuisance function $\circ$, e.g., $r_{\mu_{S_{t}}^{E}}(n)$ correspondingly to $\mu_{S_{t}}^{E}$, and further under Assumption 5.2, we have
$\hat{\omega}_{t}(\mathbf{x})-\omega_{t}(\mathbf{x})=O_{p}\left(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}}\right).$ (28)
Proof.
The lemma is immediately proved by the form of $\omega_{t}(\mathbf{x})$ as $\omega_{t}(\mathbf{x})=\mu_{S_{t}}^{E}(1,\mathbf{x})-\mu_{S_{t}}^{E}(0,\mathbf{x})+\mu_{S_{t}}^{O}(0,\mathbf{x})-\mu_{S_{t}}^{O}(1,\mathbf{x})$. ∎
We now restate our Lemma 5.4 as follows and formally prove that.
Lemma D.2.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2 and 5.3 hold, then we have
$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (29)
Proof.
We apply Proposition 1 of Kennedy et al. (Kennedy, 2023), yielding that
$\hat{f}(\mathbf{x})-f(\mathbf{x})=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+(\tilde{f}(\mathbf{x})-f(\mathbf{x}))$ (30)
$=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})|\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})|\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$. Here, $f$ is a time-series model ($1$-order autoregressive), optimized with $T-1$-length steps and $n$ samples, as shown in Eq. (8). According to Proposition 4.3, $\forall t$, $f(\mathbf{x})=\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}$. We first prove the rate of $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}$ for a fixed $t$ as follows.
$(\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})})^{2}$ (31)
$=$ $\left(\frac{\hat{\omega}_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$=$ $\left(\frac{\hat{\omega}_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})+\omega_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$=$ $\left(\frac{(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))\omega_{t}(\mathbf{x})+\omega_{t+1}(\mathbf{x})(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$\overset{(a)}{\leq}$ $\frac{2(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}\omega_{t}^{2}(\mathbf{x})+\omega_{t+1}^{2}(\mathbf{x})(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2}}{\omega_{t}^{2}(\mathbf{x})\hat{\omega}_{t}^{2}(\mathbf{x})}$
$=$ $\frac{2}{\hat{\omega}_{t}^{2}(\mathbf{x})}(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}+\frac{2\omega_{t+1}^{2}(\mathbf{x})}{\omega_{t}^{2}(\mathbf{x})\hat{\omega}_{t}^{2}(\mathbf{x})}(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2}$
$\overset{(b)}{\asymp}$ $(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}+(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2};$
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and $(b)$ is based on Assumption 5.3. Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\hat{\omega}_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds. Then, applying Lemma D.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$. Then, for $\hat{r}(\mathbf{x})$, we have the effective sample is of size $(T-1)n$, and then combining with Eq. (30) and $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$, we obtain the desired result. ∎
Appendix E Proof of Theorem 5.5
We first restate Theorem 5.5 as follows:
Theorem E.1.
Suppose Lemma 5.4 hold, then we have
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})=O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$ (32)
Proof.
As stated in Lemma 5.4, we have
$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right),$ (33)
and thus
$\hat{f}^{\mu}(\mathbf{x})-f^{\mu}(\mathbf{x})=O_{p}\left((\frac{\mu}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{\mu}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{\mu}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (34)
Under Assumption 5.2, we have
$\hat{\mu}_{Y}^{O}(a,\mathbf{x})-\mu_{Y}^{O}(a,\mathbf{x})=O_{p}(n^{-\frac{\gamma}{2\gamma+d}})$ (35)
and
$\hat{\mu}_{S_{T}}^{E}(a,\mathbf{x})-\mu_{S_{T}}^{E}(a,\mathbf{x})=O_{p}(n^{-\frac{\alpha}{2\alpha+d}})$ (36)
and
$\hat{\mu}_{S_{T}}^{O}(a,\mathbf{x})-\mu_{S_{T}}^{O}(a,\mathbf{x})=O_{p}(n^{-\frac{\beta}{2\beta+d}}).$ (37)
Then, let $a:=f^{\mu}(\mathbf{X})$,$\hat{a}=\hat{f}^{\mu}(\mathbf{X})$, $b=\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and similarly for $\hat{b}$, and under Assumption 5.3, $a$, $b$, $\hat{a}$ and $\hat{b}$ are all bounded. We analyze the term
$(ab-\hat{a}\hat{b})^{2}=$ $(ab-a\hat{b}+a\hat{b}-\hat{a}\hat{b})^{2}$ (38)
$=$ $(a(b-\hat{b})+(a-\hat{a})\hat{b})^{2}$
$\overset{(a)}{\leq}$ $2a^{2}(b-\hat{b})^{2}+2\hat{b}^{2}(a-\hat{a})^{2}$
$\overset{(a)}{\asymp}$ $(b-\hat{b})^{2}+(a-\hat{a})^{2},$
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and $(b)$ is based on the boundedness assumption. Then we have
$ab-\hat{a}\hat{b}=O_{p}\left((\frac{\mu}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{\mu}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{\mu}{(T-1)n})^{\frac{\beta}{2\beta+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}\right).$ (39)
Hence, the result is immediately proved by the form of $\tau(x)$ as $\tau(x)=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right)$ and the result of Lemma 5.4. ∎